OpenStudy (anonymous):

A torque of 0.97Nm is applied to a bicycle wheel of radius 45cm and mass 0.60kg. Treating the wheel as a hoop, find its angular acceleration.

OpenStudy (anonymous):

Have tried:\[T=1/2mr^{2}\alpha\]

OpenStudy (anonymous):

torque= moment of inertia * angular acceleration T=0.98Nm r=45cm=0.45m m=.60Kg I=1/2mr^2=0.06075Kgm^2 a=T/I=16.13168ms^-2:)

OpenStudy (anonymous):

It needs to be in radians/s^{2}

OpenStudy (anonymous):

ya that i nt imp here:)

OpenStudy (anonymous):

That is not correct.

OpenStudy (anonymous):

it means the answer nt the unit:)

OpenStudy (anonymous):

I dont understand.

OpenStudy (anonymous):

Yo

OpenStudy (anonymous):

Your answer wasnt right. I did about the same thing and still couldnt get it.

OpenStudy (anonymous):

@halipearce the answer is right............i garuntee:)

OpenStudy (anonymous):

No, I just plugged it into mastering physics. It said it was wrong.

OpenStudy (anonymous):

\[\tau = I \alpha\] I = mr^2

OpenStudy (anonymous):

Why is I=mr^2 when you are suppossed to treat the wheel like a hoop?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

moment for a hoop about an axis through its center is mr^2

OpenStudy (anonymous):

or through its center perpendicular to the plane of the hoop I should say...

OpenStudy (anonymous):

aka a wheel.